Home‎ > ‎Crew 136‎ > ‎

226-GreenHabReport

GreenHab Report
02/26/2014
Benoit Pairet

Again, the multi-plug was off, that begins to be strange. I should spend part of the night in the observatory, I'll check time to time if it's still on. I did not catch any rodents, I hope to get one tonight.

I did not water the hydroponic plants today to avoid water saturation, there is still one wellspring that does not work. I performed the transplant and it went well. I found the roots of the dead plants, I replanted them and put them under the LED lamps. I have little hope but we never know, they might grow again.


Greenhab readings for 5PM today are as follows:
                         Current             Min                 Max
Temperature              72 F            43 F            104 F
Humidity                   16 %             9 %                20 %

Greenhab readings at 9am today were as follows:

Meter           no data
Duration       no data